OpenStudy (anonymous):

quick question for everyone! What two numbers add up to 1, but when multiplied equal -3?

OpenStudy (anonymous):

*id also like to know how you solved it

OpenStudy (anonymous):

we can start with two variables \(x\) and \(y\)

OpenStudy (anonymous):

they add to 1 so \(x+y=1\) they multiply to give \(-3\) so \(xy=-3\)

OpenStudy (anonymous):

to make one variable out of this mess, \(x+y=1\) is the same as \(y=1-x\) put that in the second equation and get \[x(1-x)=-3\]

OpenStudy (anonymous):

then solve that one

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

nope haha ive been out of school for a long time, i only come here to help friends usually

OpenStudy (anonymous):

\[x(1-x)=-3\\ x-x^2=-3\\ x^2-x-3=0\] then use the quadratic formula to solve it

OpenStudy (anonymous):

im working it for now

OpenStudy (anonymous):

2.30277564, -1.30277564. it's of by 0.00000001 but its the right answer

OpenStudy (anonymous):

off* not of